MIRR A firm is considering two mutually exclusive projects, X and Y, with the following cash...

MIRR

A firm is considering two mutually exclusive projects, X and Y, with the following cash flows:

0 1 2 3 4
Project X $(1,000) $100 $320 $370 $750
Project Y $(1,000) $1,000 $110 $50 $55

The projects are equally risky, and their WACC is 13%. What is the MIRR of the project that maximizes shareholder value? Round your answer to two decimal places. Do not round your intermediate calculations.

%
Answer #1

X:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=100/1.13+320/1.13^2+370/1.13^3+750/1.13^4

=1055.52

NPV=Present value of inflows-Present value of outflows

=1055.52-1000

=$55.52(Approx)

Y:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=1000/1.13+110/1.13^2+50/1.13^3+55/1.13^4

=1039.49

NPV=Present value of inflows-Present value of outflows

=1039.49-1000

=$39.49(Approx)

Hence X is better having higher NPV.

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Future value of inflows=100*(1.13)^3+320*(1.13)^2+370*(1.13)+750

=1720.9977

MIRR=[Future value of inflows/Present value of outflows]^(1/time period)-1

=[1720.9977/1000]^(1/4)-1

=14.54%(Approx)